OpenStudy (kainui):

Here's a fun problem I just came up with using Lambert's product log W(x):

OpenStudy (kainui):

\[\LARGE W(e^x*b^b )=b*b^b\] Solve for b. Good luck! Medal to the winner. =P @ganeshie8 @dan815 @ikram002p @iambatman

OpenStudy (rational):

\(\large e^x*b^b = b*b^b e^{b*b^b}\) \(\large b^b[e^x - b e^{b*b^b}] = 0\)

OpenStudy (kainui):

You're on the right track, but you haven't solved for b yet @rational

OpenStudy (anonymous):

What is W ? a constant ?

OpenStudy (kainui):

@HatcrewS nope, it's the product log function, check it out: https://luckytoilet.wordpress.com/tag/product-log/ http://www.had2know.com/academics/lambert-w-function-calculator.html

OpenStudy (anonymous):

then what is the value of W

OpenStudy (kainui):

it has no value, just like "ln" has no value. It's a function.

OpenStudy (anonymous):

why is the the exponent of the left hand side equal to the whole left hand side.

OpenStudy (anonymous):

makes no sense @@

OpenStudy (kainui):

What do you mean? I don't follow

OpenStudy (anonymous):

it says that when u remove an X from the exponent, the power doesnt change

OpenStudy (kainui):

where, show me the equation you're talking about

OpenStudy (kainui):

At any rate, continuing where rational left off we can conclude from \[\LARGE b^b(e^x−be^{bb^b})=0\] this means either \[\LARGE b^b=0\] or \[\LARGE e^x−be^{bb^b}=0\] The first one is pretty obviously \[\LARGE b \ln b = \ln0=-\infty\] but that's not really the answer we're looking for, since it's not a number. So let's try to continue with the second one to get the real equation for b in terms of x.

OpenStudy (rational):

wolfram doesn't like either of the equations

OpenStudy (kainui):

Hahaha well I'm sorry you can't cheat.

OpenStudy (rational):

\(\large e^x - b e^{b*b^b}= 0 \implies x = \ln b + b*b^b\)